please answer all questions and show all steps

3. Draw the magnitude characteristic of the Bode plot of the following transfer function: Hsout20(0.1s + 1 Vin S(s1)(0.01s +1) a) ldentify the poles and zeros b) Sketch the magnitude plot. The y-axis should be in dB. The x-axis should logarithmic, but labeled in rad/sec. Use MATLAB to draw the complete Bode plot and check your answer. If the input current is uln(t)-2 cos(2t + 30%), what is the output voltage. Use your MATLAB Bode plot to answer this. c) d)
Answer #1

Ans)

a) from the given transfer function putting s=jw

20(0.1jw 1) H(jw) =

H(jw)=\frac{20(j\frac{w}{10}+1)}{jw(j\frac{w}{1}+1)(j\frac{w}{100}+1)}

poles are roots of denominator polynomial i.e

s=0,s=-1,s=-100  

poles\: \: at\: \: \: w=0\: \: rad/s,\: \: w=1\: \: rad/s,\: \: w=100\: \: rad/s

zeroes are roots of numerator

s=-10

zero\: \: at\: \: \: w=10\: \: rad/s

===================

b)

For magnitude plot we find asymptodes first comes pole at w=0\: \: rad/s contributes -20dB/decade slope gain at w=0.1 rad/sec is

G_{w=0.1}=20log_{10}\frac{20}{w}=20log_{10}\frac{20}{0.1}=46.02\: \: dB

Next comes pole at w=1 rad/s contributes -20dB decade net slope is

-20dB/decade-20dB/decade=-40dB/decade

Next comes zero\: \: at\: \: \: w=10\: \: rad/s it contributes slope of +20dB/decade ,now net slope of the line to be drawn is -40dB/decade+20dB/decade=-20dB/decade

Next comes pole at w=100 rad/s contributes -20dB decade net slope is

-20dB/decade-20dB/decade=-40dB/decade

using the above details magnitude plot is

====================

c)

The matlab code for bode plot given below ,convolution command used for multiplication of polynomials in the denominator of transferfunction

================

d)

Given input v_{IN}(t)=2cos(2t+30^{o})\: \: V from this w=2\: \: rad/s the gain and phase at w=2 rad/s is marked in the bode plot above

Gain(dB)=12.9\: \: dB

Gain=10^{\frac{12.9}{20}}=4.416

Gain=4.416

phase=-144^{o} from plot using this output voltage is

v_{out}(t)=gain*v_{IN}(t)\angle -144^{o}=4.416*2cos(2t+30^{o}-144^{o})\: \: V

\mathbf{v_{out}(t)=8.83cos(2t-114^{o})\: \: V}